OpenStudy (anonymous):

Describe the meaning of the rate. 6 days each month A. Something happens 6 days per month. B. Each month is 6 days long. C. Each day is one month long. D. A day happens six times each month.

OpenStudy (anonymous):

im thinking its A but which do you think it is?

OpenStudy (anonymous):

D.

OpenStudy (anonymous):

@vampiregodess Can You Explain this To Me ??

OpenStudy (anonymous):

ok 6 days each month is making it sound like somthing happens 6 days out of the month ex: six days each month i go swimming, six days each month it rains ect ect

OpenStudy (anonymous):

OHHHHHHH

OpenStudy (anonymous):

Will You Help Me With 4 More @vampiregodess

OpenStudy (anonymous):

yes i will try my best to help you :)

OpenStudy (anonymous):

Thank You So Much !

OpenStudy (anonymous):

An airplane travels at an average speed of 250 mph for 3 hours. How far does the airplane travel? A. 83.3 mi B. 250 mi C. 500 mi D. 750 mi

OpenStudy (anonymous):

ok the question is asking how many miles it travels in 3 hours and it goes about 250 miles per a hour all you have to do is 250+250+250=?

OpenStudy (anonymous):

So D.750 ?

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

The Peytons' car averaged 66 miles per hour on their trip. The trip took 90 minutes. How far did they travel? A. 90 mi B. 99 mi C. 156 mi D. 5940 mi @vampiregodess

OpenStudy (anonymous):

ok this one is a little harder they traveld 1 hour and 30 min so its 66+(\[66 \div 2\])

OpenStudy (anonymous):

So Uhh 99 ??

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

brb im going to go and eat lunch

OpenStudy (anonymous):

@Imperfect_Queen im back

OpenStudy (anonymous):

Javi figures that 2 small pizzas will feed 3 people. He expects 15 people at the team pizza party. How many small pizzas will he need for 15 people? A. 6 B. 10 C. 30 D. 45

OpenStudy (anonymous):

Last One

OpenStudy (anonymous):

@vampiregodess

OpenStudy (anonymous):

5+5=?

OpenStudy (anonymous):

@Imperfect_Queen

OpenStudy (anonymous):

its B 15/3=5 5+5=10

OpenStudy (anonymous):

TY